Can you please help me graph this linear inequalities y<6x+2
Answer by mananth(16946)

y>6x+2
First determine the boundary of the inequality
The boundary
y=6x+2
when x=0 y=2 (0,2)
when y=0, x=-2/3 =-1/3 (-1/3,0)

Since y > than something the upper side of the line will be shaded. The line will be dotted


m.ananth@hotmail.ca